On May 5, 1961, Alan Shepard, Jr., became the first American to travel into outer space. His Mercury space capsule was carried on top of a Redstone rocket, which was a converted Intercontinental Ballistic Missile that propelled his capsule into a 15-minute sub-orbital flight with a maximum altitude of 117 miles and a maximum speed of 5,134 mph. During the boost phase of the flight, Commander Shepard experienced acceleration that was between 5 and 6 G’s .

One "G" of acceleration is the rate at which an object will accelerate if it is dropped near the Earth's surface, ignoring air resistance. One "G" of force is the force necessary to generate one G of acceleration (although "G" is not an official SI unit of force).

Suppose you built a rocket-powered dragster which could produce a constant acceleration, for a few seconds, that was similar to the acceleration that Cmdr Shepard experienced for the first several minutes of his flight.

On one pass, your dragster produces a constant acceleration of exactly 5.39 G's throughout the whole race.

How fast would your dragster be going at the end of a drag strip that was exactly 1000 feet long?

Your answer must be in miles/hour.

Hint: You do not need to know or estimate the race’s elapsed time to find the answer.
Answer #1

a = 5.39 G's = 5.39 * 9.8 = 52.822 m/s2

d = 1000 feet = 304.8 m

Assuming that dragster starts from rest,

v2 - u2 = 2ad

v = sqrt ( 2ad)

v = 179.44 m/s

1 m/s = 2.24 mph

so,

v = 401.95 mph